The diagonal of a rectangle is 13cm. How long is its width if its length is 10cm?
Answer by vidya p(12)

Solution :
Let ABCD is an rectangle ,
we know that all angles of an rectangle is an right angle,
Hence, the diagonal divides the rectangle in two right angled triangle.
Let take one right angled triangle, ABC
in which Angle B = 90 degree,
and diagonal AC = 13cm and AB = 10 cm
hence by using pythagorus theorem , we get :
Ac^2 = AB^2 + BC^2
13^2 = 10^2 + BC^2
BC^2 = 13^2 - 10^2
BC^2 = ( 13 - 10 ) ( 13 + 10)
BC^2 = 3 * 23
BC = sqrt ( 69)
Thanks ,
Best regards.